OpenStudy (anonymous):

Which ordered pair is a solution of this equation? -3x - 7y=8 can someone explain how to plug in thse numbers please and thank you

OpenStudy (bibby):

Do you mean how to check whether or not an ordered pair is a solution? Just plug the x and y in where (x, y)

OpenStudy (anonymous):

do i multiply subtract ?

OpenStudy (bibby):

For example, is (3,4) a solution? Plug in x=3 y=4

OpenStudy (anonymous):

ya i know it's going to lok like this -3(3)-7(-2)= 8 but do i subtract multiply the terms or

OpenStudy (bibby):

PEDMAS Distribute the -3 and the 7 and then subtract

OpenStudy (anonymous):

-6-9=-15

OpenStudy (bibby):

not quite -3(3)-7(-2)=[-3*3]-[7*-2]

OpenStudy (anonymous):

so out of these answers which would be correct? 3,-2 -2,2 -2,3 2,-2

OpenStudy (bibby):

set x = to 0 and solve for y